# Causal Tracing of Object Representations in Large Vision Language Models: Mechanistic Interpretability and Hallucination Mitigation

### v2

@@ Introduction
Large Vision-Language Models (LVLMs) have rapidly evolved, demonstrating impressive capabilities across diverse tasks. However, existing interpretability studies fall short in capturing the full complexity of visual information flow, thereby limiting progress in critical downstream applications such as hallucination mitigation. Fundamental questions require further investigation, particularly regarding how LVLMs process visual object features and align them with textual semantics in cross-modal representations. Furthermore, how visual and textual tokens elicit distinct functional behaviors from three core model components—such as multi-head self-attention (MHSA), feed-forward networks (FFNs), and hidden states—across layers, remains underexplored. Previous studies have partially addressed these questions but still leave notable limitations. Attention knockout experiments (neo2024towards) demonstrate that LVLMs extract object information from visual object tokens in the middle to late layers. However, it lacks an analysis of the cross-modal interactions between visual and textual tokens, as well as the functional roles of the MLP and hidden states. NOTICE (golovanevsky2024vlms) introduces semantic image pairs for image corruption and symmetric token replacement for text corruption to analyze how MHSA and MLP contribute to information aggregation in textual tokens, but overlooks the effect of visual tokens.
@@
FCCT not only reveals cross-modal information flow of visual objects, but also provides valuable guidance for hallucination mitigation. Prior studies (tang2025mitigating) suggest that deep unidirectional information flow can lead to the progressive degradation of fine-grained semantic cues encoded in earlier layers, which may be a potential cause of object hallucination. To prevent mid-layer degradation of critical information during forward and reinforce components with strong causal effects identified by FCCT, we further propose Intermediate Representation Injection (IRI), a training-free inference-time technique that injects crucial mid-layer representations into subsequent layers, thereby enhancing visual perception capability and mitigating hallucination. FCCT offers fine-grained and quantitative guidance for selecting model components and layers, serving as a theoretical foundation for the design and implementation of the IRI method. Consistent improvement across five widely used benchmarks and five advanced LVLMs demonstrates that IRI achieves state-of-the-art (SOTA) performance.
@@
While LVLMs have demonstrated remarkable capabilities across various downstream tasks, their mechanistic interpretability remains underexplored. Existing interpretability methods, such as probing (salin2022vision), activation patching (basu2024understanding; palit2023towards; golovanevsky2024vlms), logit lens (neo2024towards; huo2024mmneuron) often provide only a coarse-grained analysis of model components or do not fully disentangle the complex interactions between visual and textual token representations. In contrast, our work employs causal tracing with Gaussian noise to precisely quantify the functional roles of MHSA, FFN, and hidden states for both visual and textual tokens across layers, enabling a fine-grained analysis of how LVLMs perceive and process visual object information.
@@
LVLMs frequently produce content that deviates from visual information, leading to object hallucination. Existing hallucination mitigating strategies can be broadly categorized into three types: (1) Training-based approaches enhance model factuality by pre-training or finetuning with carefully curated datasets (yu2024rlaif; you2023ferret) and novel training objectives (lyu2024alleviating). These methods can be effective but require substantial annotation and computational resources. (2) Contrastive decoding (leng2024mitigating; huang2024opera; zhong2024investigatingmitigatingmultimodalhallucination) leverages differences between deliberately perturbed decoding paths to promote generations that are more consistent in visual information. However, such methods introduce significant latency at inference time. (3) Inference-time interventions modify internal activations such as attention heads outputs (liu2024paying; li2025cai) or hidden states (liu2024reducing) to steer the model toward more faithful outputs. However, these methods generally lack interpretability regarding the selection of layers and components.
@@
We restrict our scope to LVLMs that are based on auto-regressive Transformer architecture (vaswani2017attention), as it is adopted by most SOTA LVLMs. The model receives as input a visual input sequence $\boldsymbol{V}=\{v_{1},v_{2},\dots,v_{m}\}$ and a textual input sequence $\boldsymbol{T}=\{t_{1},t_{2},\dots,t_{n}\}$, where $m$ and $n$ denote the sequence lengths of the visual and textual input. The textual and visual input sequences are concatenated together and processed through $L$ transformer layers of the language decoder, each consisting of multi-head self-attention (MHSA), feed-forward network (FFN) that is usually a multilayer perception (MLP), and a residual stream is applied between each components. The $l$-th layer hidden state $\boldsymbol{h}^{\left(l\right)}$ can be computed from the previous layer:
@@
Causal tracing (also known as activation patching or causal mediation analysis) is a widely used interpretability technique that selectively replaces internal activations to probe the causal contribution of specific model components (meng2022locating). In the context of large language models (LLMs), causal tracing is frequently employed to examine the storage and retrieval mechanisms of factual associations (meng2022locating), and document-level relevance (liu2025large). In the context of LVLMs, we are the first to propose using controlled Gaussian noise perturbations on input images for causal tracing. By adding controlled Gaussian noise to the entire image and then selectively restoring the activations of specific components, we conduct a fine-grained analysis of the internal mechanisms responsible for visual object perception and comprehension in LVLMs.
Specifically, we select 500 images from the COCO dataset and design object-related questions for the objects present in each image. Following the analysis methodology of ROME (meng2022locating), we define three types of inference runs:
@@
For each attention source layer $k\in\mathcal{L}_{\text{src}}^{\text{attn}}$, we record the MHSA output of the last token: $\boldsymbol{a}^{(k)}\in\mathbb{R}^{d}$. For each target layer $l\in\mathcal{L}_{\text{tgt}}^{\text{attn}}$ such that $l>k$, we inject the stored activations into the target MHSA output, and scaling them by RRRR to modulate their contribution to reflect causal effect:
@@
We adopt the widely used LLaVA-1.5-7b (liu2024improved), Qwen-VL-Chat (bai2023qwen), LLaVA-NeXT (liu2024llavanext), Qwen2-VL-7B(wang2024qwen2) and InternVL2-8B (chen2024far) as baseline LVLMs.
@@
  • POPE (li2023evaluating) employs a binary question-answering format, inquiring LVLMs to answer if a special object exists in the given image. Following previous works, we adopt Accuracy and F1 score as the metrics.
  • MME (Fu2023MMEAC) serves as a comprehensive tool for assessing the capabilities of LVLMs across both 10 perception tasks and 4 cognition tasks. Consequently, task scores are reported as the evaluation metric.
  • CHAIR (rohrbach2018object) is a widely used metric for assessing object hallucination in responses of LVLMs. The CHAIR metric comprises two important indicators, denoted as ${C}_{S}$ and ${C}_{I}$, with the following calculation formulas:
@@
  • MMHal-Bench (sun2023aligning) comprises 96 meticulously designed questions, which evaluates response-level hallucination rate (VH.%) and informativeness (Score). It asks GPT-4 to compare model outputs with human responses and object labels for evaluation.
  • MHumanEval (yu2024rlaif) is designed to evaluate hallucination performance by human annotators. The benchmark contains 146 samples collected from Object HalBench and MMHal-Bench. Given model responses, we ask three human annotators to label the hallucinated segments and compute the mean response-level hallucination rate (Hu.%) as the evaluation metric.
@@
We compared our proposed IRI method with the following SOTA training-free methods: VCD (leng2024mitigating) contrasts model logits derived from original and distorted visual input to reduce the over-reliance on statistical bias and unimodal priors. OPERA (huang2024opera) introduces a penalty term on the model logits during the beam-search decoding to mitigate the over-trust issue. PAI (liu2024paying) intervenes on attention heads by leveraging their original direction and optimizes the output distribution during decoding to mitigate language bias. VTI (liu2024reducing) mitigates hallucination by steering layer-wise hidden states during inference to enhance visual feature stability.

### v3

@@ Introduction
Large Vision-Language Models (LVLMs) have rapidly evolved, demonstrating impressive capabilities across diverse tasks. However, existing interpretability studies fall short in capturing the full complexity of visual information flow, thereby limiting progress in critical downstream applications such as hallucination mitigation. Fundamental questions require further investigation, particularly regarding how LVLMs process visual object features and align them with textual semantics in cross-modal representations. Furthermore, how visual and textual tokens elicit distinct functional behaviors from three core model components—such as multi-head self-attention (MHSA), feed-forward networks (FFNs), and hidden states—across layers, remains underexplored. Previous studies have partially addressed these questions but still leave notable limitations. Attention knockout experiments (Neo et al. 2024) demonstrate that LVLMs extract object information from visual object tokens in the middle to late layers. However, it lacks an analysis of the cross-modal interactions between visual and textual tokens, as well as the functional roles of the MLP and hidden states. NOTICE (Golovanevsky et al. 2024) introduces semantic image pairs for image corruption and symmetric token replacement for text corruption to analyze how MHSA and MLP contribute to information aggregation in textual tokens, but overlooks the effect of visual tokens.
@@
FCCT not only reveals cross-modal information flow of visual objects, but also provides valuable guidance for hallucination mitigation. Prior studies (Tang et al. 2025) suggest that deep unidirectional information flow can lead to the progressive degradation of fine-grained semantic cues encoded in earlier layers, which may be a potential cause of object hallucination. To prevent mid-layer degradation of critical information during forward and reinforce components with strong causal effects identified by FCCT, we further propose Intermediate Representation Injection (IRI), a training-free inference-time technique that injects crucial mid-layer representations into subsequent layers, thereby enhancing visual perception capability and mitigating hallucination. FCCT offers fine-grained and quantitative guidance for selecting model components and layers, serving as a theoretical foundation for the design and implementation of the IRI method. Consistent improvement across five widely used benchmarks and five advanced LVLMs demonstrates that IRI achieves state-of-the-art (SOTA) performance.
@@
While LVLMs have demonstrated remarkable capabilities across various downstream tasks, their mechanistic interpretability remains underexplored. Existing interpretability methods, such as probing (Salin et al. 2022), activation patching (Basu et al. 2024; Palit et al. 2023; Golovanevsky et al. 2024), logit lens (Neo et al. 2024; Huo et al. 2024), in-context learning (Li et al. 2025d, b, c) provide only a coarse-grained analysis of model components or do not fully disentangle the complex interactions between visual and textual token representations. In contrast, our work employs causal tracing with Gaussian noise to precisely quantify the functional roles of MHSA, FFN, and hidden states for both visual and textual tokens across layers, enabling a fine-grained analysis of how LVLMs perceive and process visual object information.
@@
LVLMs frequently produce content that deviates from visual information, leading to object hallucination. Existing hallucination mitigating strategies can be broadly categorized into three types: (1) Training-based approaches enhance model factuality by pre-training or finetuning with carefully curated datasets (Yu et al. 2024; You et al. 2023; Zhang et al. 2025) and novel training objectives (Lyu et al. 2024). These methods can be effective but require substantial data and computational resources. (2) Contrastive decoding (Leng et al. 2024; Huang et al. 2024; Zhong et al. 2024) leverages differences between deliberately perturbed decoding paths to promote generations that are more consistent in visual information. However, such methods introduce significant latency at inference time. (3) Inference-time interventions modify internal activations such as attention heads outputs (Liu, Zheng, and Chen 2024; Li et al. 2025a; Ye et al. 2025) or hidden states (Liu, Ye, and Zou 2024) to steer the model toward more faithful outputs. However, these methods generally lack interpretability of the selection of layers and components.
@@
We restrict our scope to LVLMs that are based on auto-regressive Transformer architecture (Vaswani et al. 2017), as it is adopted by most SOTA LVLMs. The model receives as input a visual input sequence $\boldsymbol{V}=\{v_{1},v_{2},\dots,v_{m}\}$ and a textual input sequence $\boldsymbol{T}=\{t_{1},t_{2},\dots,t_{n}\}$, where $m$ and $n$ denote the sequence lengths of the visual and textual input. The textual and visual input sequences are concatenated together and processed through $L$ transformer layers of the language decoder, each consisting of multi-head self-attention (MHSA), feed-forward network (FFN) that is usually a multilayer perception (MLP), and a residual stream is applied between each components. The $l$-th layer hidden state $\boldsymbol{h}^{\left(l\right)}$ can be computed from the previous layer:
@@
Causal tracing (also known as activation patching or causal mediation analysis) is a widely used interpretability technique that selectively replaces internal activations to probe the causal contribution of specific model components (Meng et al. 2022). In the context of large language models (LLMs), causal tracing is frequently employed to examine the storage and retrieval mechanisms of factual associations (Meng et al. 2022), and document-level relevance (Liu, Mao, and Wen 2025). In the context of LVLMs, we are the first to propose using controlled Gaussian noise perturbations on input images for causal tracing. By adding controlled Gaussian noise to the entire image and then selectively restoring the activations of specific components, we conduct a fine-grained analysis of the internal mechanisms responsible for visual object perception and comprehension in LVLMs.
Specifically, we select 500 images from the COCO dataset and design object-related questions for the objects present in each image. Following the analysis methodology of ROME (Meng et al. 2022), we define three types of inference runs:
@@
For each attention source layer $k\in\mathcal{L}_{\text{src}}^{\text{attn}}$, we record the MHSA output of the last token: $\boldsymbol{a}^{(k)}\in\mathbb{R}^{d}$. For each target layer $l\in\mathcal{L}_{\text{tgt}}^{\text{attn}}$ such that $l>k$, we inject the stored activations into the target MHSA output, and scaling them by $RR$ to modulate their contribution to reflect causal effect:
@@
We adopt the widely used LLaVA-1.5-7b (Liu et al. 2024a), Qwen-VL-Chat (Bai et al. 2023), LLaVA-NeXT (Liu et al. 2024b), Qwen2-VL-7B(Wang et al. 2024) and InternVL2-8B (Chen et al. 2024) as baseline LVLMs.
@@
  • POPE (Li et al. 2023) employs a binary question-answering format, inquiring LVLMs to answer if a special object exists in the given image. Following previous works, we adopt Accuracy and F1 score as the metrics.
  • MME (Fu et al. 2023) serves as a comprehensive tool for assessing the capabilities of LVLMs across both 10 perception tasks and 4 cognition tasks. Consequently, task scores are reported as the evaluation metric.
  • CHAIR (Rohrbach et al. 2018) is a widely used metric for assessing object hallucination in responses of LVLMs. The CHAIR metric comprises two important indicators, denoted as ${C}_{S}$ and ${C}_{I}$, with the following calculation formulas:
@@
  • MMHal-Bench (Sun et al. 2023) comprises 96 meticulously designed questions, which evaluates response-level hallucination rate (VH.%) and informativeness (Score). It asks GPT-4 to compare model outputs with human responses and object labels for evaluation.
  • MHumanEval (Yu et al. 2024) is designed to evaluate hallucination performance by human annotators. The benchmark contains 146 samples collected from Object HalBench and MMHal-Bench. Given model responses, we ask three human annotators to label the hallucinated segments and compute the mean response-level hallucination rate (Hu.%) as the evaluation metric.
@@
We compared our proposed IRI method with the following SOTA training-free methods: VCD (Leng et al. 2024) contrasts model logits derived from original and distorted visual input to reduce the over-reliance on statistical bias and unimodal priors. OPERA (Huang et al. 2024) introduces a penalty term on the model logits during the beam-search decoding to mitigate the over-trust issue. PAI (Liu, Zheng, and Chen 2024) intervenes on attention heads by leveraging their original direction and optimizes the output distribution during decoding to mitigate language bias. VTI (Liu, Ye, and Zou 2024) mitigates hallucination by steering layer-wise hidden states during inference to enhance visual feature stability.
@@ Acknowledgement
Xiaocheng Feng and Xiachong Feng are the co-corresponding authors of this work. We thank the anonymous reviewers for their insightful comments. This work was supported by the National Natural Science Foundation of China (NSFC) (grant 62522603, 62276078, U22B2059), the Key R&D Program of Heilongjiang via grant 2022ZX01A32, and the Fundamental Research Funds for the Central Universities ( XNJKKGYDJ2024013 ).